OpenStudy (anonymous):

Okay.. now time for multiplying fractions. ): 5/2 x 16/5? Someone please explain... Do I still make the demoniaters equal?

OpenStudy (anonymous):

cancel 5 and you will get 16/2 which is 8

OpenStudy (anonymous):

how do you get 8 from 16/2?

OpenStudy (anonymous):

(5/2) * (16/5) = 80/10 = 8

OpenStudy (anonymous):

ahh! Okay that makes sense kindof .

OpenStudy (anonymous):

(5*16)/(2*5)

OpenStudy (amistre64):

\[\frac 52 *\frac {16}5\] \[\frac {16}2 *\frac {5}5\] \[\frac {16}2 *1\] \[\frac {16}2=8\]

OpenStudy (anonymous):

you don't need to as it is multiplication and not addition or subtraction. you can cancel 5 in the numerator with the 5 in the denominator. after this you will be left with 16/2 which is 16 divided by 2. Since 2*8=16 therefore 16/2=8.

OpenStudy (anonymous):

alright that makes sense. But what about division? like.. 6/5 divided by 3/5? can you cancel out the 5s again?

OpenStudy (amistre64):

you can cancel out the 1/5s

OpenStudy (anonymous):

6/5 divided by 3/5 can be written as (6/5)*(5/3). therefore you can cancel the 5's and the answer will be 6/3=2.

OpenStudy (amistre64):

\[\frac{6*\cancel{\frac15}}{3*\cancel{\frac 15}}=\frac 63\]

OpenStudy (anonymous):

alright, thanks! It makes sense.